# DeTrigger: A Gradient-Centric Approach to Backdoor Attack Mitigation in Federated Learning

### v1

@@ 1. Introduction
Federated Learning (FL) is a decentralized machine learning approach that trains a global model by aggregating locally trained models from mobile and embedded devices (McMahan et al., 2017; Shin et al., 2024). This method leverages distributed data and computational resources, reducing the dependency on centralized processing (Park et al., 2023; Yao et al., 2021; Park and Ko, 2024). Federated learning powers mobile applications, such as sensor data analysis (Park et al., 2020; Shen et al., 2024; Ouyang et al., 2022), autonomous vehicle (Li et al., 2021c; Pokhrel and Choi, 2020), and real-time computer vision (Ahn et al., 2023; Ouyang et al., 2023; Deng et al., 2023; Yun et al., 2024), by using large, diverse datasets without data sharing. A key principle is preserving local data privacy, as the server aggregates updates without accessing raw data (McMahan et al., 2017; Li et al., 2021b; T Dinh et al., 2020). However, this also means the server cannot verify updates, making federated learning vulnerable to model-poisoning attacks from malicious clients (Bagdasaryan et al., 2020; Fang et al., 2020; Cao and Gong, 2022).
@@
Our proposed mechanism, DeTrigger, leverages gradient analysis techniques for backdoor trigger detection. By examining model gradients, which capture how model weights respond to varying inputs, we can detect subtle deviations indicative of a backdoor trigger. However, this approach presents challenges in distinguishing backdoor attack components from inherent input-specific noise. DeTrigger addresses this issue by isolating trigger-dependent features within its gradient preprocessing operations. This method is particularly scalable for federated learning, as gradient-based analysis enables efficient abnormal pattern detection without requiring exhaustive inspection of client models on specific triggers. Furthermore, by focusing on gradient behavior, DeTrigger minimizes the need for extensive verification datasets, which would otherwise be impractical in large-scale distributed systems.
@@ 2.1. Backdoor Attacks
A backdoor attack manipulates a model $f_{back}(\cdot)$ to produce a designated incorrect output $Y_{back}$ for input $X_{t}$ with a trigger $T$ is provided while maintaining normal behavior $Y$ on unaltered data $X$. This can be formulated as $X_{t}=(1-M)X+MT,Y=f_{back}(X),Y_{back}=f_{back}(X_{t})$ (Jin et al., 2022). Such attacks are typically executed by training the model with manipulated data that incorporates the trigger pattern. Gu et al. (Gu et al., 2017) demonstrated the vulnerability of deep neural networks to backdoor attacks by introducing pixel pattern-based triggers as a method of embedding backdoors in models.
For example, in an image-based traffic sign recognition system for autonomous vehicles, an attacker could introduce a subtle trigger, such as a small sticker, linked to a specific attack label (e.g., a speed limit sign or a stop sign). Once the model is trained with this poisoned data, it becomes susceptible to backdoor exploitation. When deployed, the model misinterprets any sign containing this trigger as the designated label, potentially causing the vehicle to make dangerous decisions, leading to traffic violations, accidents, or life-threatening situations. This example highlights the serious risks backdoor attacks pose in safety-critical systems like autonomous driving (Li et al., 2021d; Wenger et al., 2021).
@@
Alternatively, anomaly detection techniques focus on identifying abnormal patterns in model updates, such as deviations in weight distributions or gradients (Yin et al., 2018; Blanchard et al., 2017). While potentially more accurate, approaches based on update pattern analysis tend to incur high computational costs and are less scalable, particularly when dealing with large models or a high number of clients. Additionally, recently proposed methods, such as trigger pattern detection and update similarity analysis, attempt to directly identify backdoor triggers by analyzing model behaviors on synthetic or auxiliary data (Elhattab et al., 2023; Wang et al., 2019; Guo et al., 2019). However, these approaches also face challenges in terms of efficiency and scalability, making the development of effective and scalable defenses against backdoor attacks in federated learning an ongoing challenge.
@@ 3. Core Idea and Feasibility Study
Figure 3. Overall workflow of DeTrigger. DeTrigger leverages insights from adversarial attack methodologies to effectively identify trigger and prune the backdoor knowledge.
To counter trigger-based backdoor attacks, identifying the trigger pattern is crucial, as tampered models act benignly without it. This section highlights our defense’s core idea, inspired by the relationship between backdoor and adversarial attacks, followed by a feasibility study.
Core Idea. Backdoor and adversarial attacks share overlapping techniques and goals. Adversarial attacks identify subtle input perturbations to alter predictions, while backdoor attacks use a predefined trigger to produce similar incorrect outputs. This raises the question: “Can adversarial techniques reverse-engineer the backdoor trigger?” While both share commonalities, they exhibit distinct differences. As shown in Figure 2 (a), adversarial attacks compute input-specific perturbations, whereas backdoor attacks apply a consistent trigger across inputs. Additionally, adversarial attacks can target any model, while backdoor attacks affect only models trained with trigger samples.
@@ 4. Threat Model and Assumptions
Assumptions. We assume that a federated learning system with numerous clients participating without directly sharing their local training data with the central server. Additionally, we assume that malicious clients comprise less than 50% of the total number of participating clients(Fang et al., 2020). We consider these assumptions reasonable for practical federated learning environments and associated attack scenarios. In addition, we assume that the server has a few clean validation samples that can be used for additional processes.
@@ 5.2. Gradient Preprocessing
Figure 5. (a) Conceptual illustration of the impact of temperature scaling on normal data feature space and backdoor feature space. (b) L1-norm between ground truth and inferred triggers with varying temperatures. (c) Sample of ground truth and inferred triggers with different temperatures.
@@
Using these input layer gradients, DeTrigger identifies and filters additive elements within each model’s gradient that shift predictions from the original to the target attack label. As also demonstrated in our preliminary studies (Figure 2 (c)), this process extracts information within the gradients specifically linked to the trigger pattern. To minimize sample-specific noise variations, DeTrigger averages the gradient elements across multiple samples. The averaged gradients are then normalized through min-max scaling, and a mask is generated to emphasize spatial locations where gradient amplitudes exceed a set threshold. This mask allows DeTrigger to focus on regions with stronger trigger-related signals, effectively filtering out irrelevant information. In this paper, we set the threshold to 0.5.
@@
This insight aligns with prior findings by Su et al., who observed that the decision boundary of the backdoor sample is tangent to other labels (Su et al., 2024). Figures 5 (b) and (c) illustrate the impact of temperature scaling on the L1-norm and inferred trigger patterns, respectively. Specifically, the results show the L1-norm of gradients decreases with moderate temperature scaling ($T>1$), improving the clarity of inferred trigger patterns. This suggests gradients obtained with smoothed probability distributions (i.e., higher temperatures) can more effectively capture information related to the backdoor trigger. Note that we set the temperature to five in this work.
@@ 8. Conclusion
Acknowledgements.
We acknowledge Sungmin Lee, Ph.D. candidate at Yonsei University, for the constructive discussion about the work.

### v2

@@ 1. Introduction
Federated Learning is a decentralized machine learning approach that trains a global model by aggregating locally trained models from mobile and embedded devices (McMahan et al., 2017; Shin et al., 2024). This method leverages distributed data and computational resources, reducing the dependency on centralized processing (Park et al., 2023; Yao et al., 2021; Park and Ko, 2024). Federated learning powers various ubiquitous computing applications, such as sensor data analysis (Park et al., 2020; Shen et al., 2024; Ouyang et al., 2022), autonomous vehicle (Li et al., 2021c; Pokhrel and Choi, 2020), and real-time computer vision (Ahn et al., 2023; Ouyang et al., 2023; Deng et al., 2023; Yun et al., 2024), by using large, diverse datasets collected from distributed mobile and embedded devices without data sharing. Since such mobile applications often exploit privacy sensitive data (Ra et al., 2017; Lee et al., 2025), a key principle of federated learning is preserving local data privacy, as the server aggregates updates without accessing raw data (McMahan et al., 2017; Li et al., 2021b; T Dinh et al., 2020). However, this also means the server cannot verify updates, making federated learning vulnerable to model-poisoning attacks from malicious clients (Bagdasaryan et al., 2020; Fang et al., 2020; Cao and Gong, 2022). Considering a ubiquitous computing environment’s distributed and collaborative nature, such malicious participants introduce severe security threats to the federated learning systems (Liu et al., 2021; Zhang et al., 2024).
@@
Our proposed mechanism, DeTrigger, leverages gradient analysis techniques which is widely used for crafting adversarial perturbation (Goodfellow, 2014; Carlini and Wagner, 2017) for backdoor trigger detection. By examining model gradients, which capture how model weights respond to varying inputs, we can detect subtle deviations indicative of a backdoor trigger. However, this approach presents challenges in distinguishing backdoor attack components from inherent input-specific noise. DeTrigger addresses this issue by isolating trigger-dependent features within its gradient preprocessing operations. This method is particularly scalable for federated learning, as gradient-based analysis enables efficient abnormal pattern detection without requiring exhaustive inspection of client models on specific triggers. Furthermore, by focusing on gradient behavior, DeTrigger minimizes the need for extensive verification datasets, which would otherwise be impractical in large-scale distributed systems.
@@ 2.1. Backdoor Attacks
A backdoor attack manipulates a model $f_{back}(\cdot)$ to produce a designated incorrect output $Y_{back}$ for input $X_{t}$ with a trigger $T$ is provided while maintaining normal behavior $Y$ on unaltered data $X$. This can be formulated as $X_{t}=(1-M)X+MT,Y=f_{back}(X),Y_{back}=f_{back}(X_{t})$ where $M$ denotes a mask to inject the trigger pattern in specific location (Jin et al., 2022). Such attacks are typically executed by training the model with manipulated data that incorporates the trigger pattern. Gu et al. (Gu et al., 2017) demonstrated the vulnerability of deep neural networks to backdoor attacks by introducing pixel pattern-based triggers as a method of embedding backdoors in models.
For example, in an image-based traffic sign recognition system for autonomous vehicles, an attacker could introduce a subtle trigger, such as a small sticker, linked to a specific attack label (e.g., a speed limit sign or a stop sign) (Chen et al., 2017). Once the model is trained with this poisoned data, it becomes susceptible to backdoor exploitation. When deployed, the model misinterprets any sign containing this trigger as the designated label, potentially causing the vehicle to make dangerous decisions, leading to traffic violations, accidents, or life-threatening situations. This example highlights the serious risks backdoor attacks pose in safety-critical systems like autonomous driving (Li et al., 2021d; Wenger et al., 2021).
@@
Alternatively, anomaly detection techniques focus on identifying abnormal patterns in model updates, such as deviations in weight distributions or gradients (Yin et al., 2018; Blanchard et al., 2017). While potentially more accurate, approaches based on update pattern analysis tend to incur high computational costs and are less scalable, particularly when dealing with large models or a high number of clients. Additionally, recently proposed methods, such as trigger pattern detection and local gradient similarity analysis, attempt to directly identify backdoor triggers by analyzing model behaviors on synthetic or auxiliary data (Elhattab et al., 2023; Wang et al., 2019; Guo et al., 2019). However, these approaches also face challenges in terms of efficiency and scalability, making the development of effective and scalable defenses against backdoor attacks in federated learning an ongoing challenge.
@@ 3. Core Idea and Feasibility Study
To counter trigger-based backdoor attacks, identifying the trigger pattern is crucial, as tampered models act benignly without it. This section highlights our defense’s core idea to identify the trigger-pattern, inspired by the relationship between the backdoor and adversarial attacks, followed by a feasibility study.
Core Idea. Backdoor and adversarial attacks share overlapping techniques and goals. Adversarial attacks identify subtle input perturbations to alter predictions, while backdoor attacks use a predefined trigger to produce similar incorrect outputs. This raises the question: “Can adversarial techniques reverse-engineer the backdoor trigger?” While both share commonalities, they exhibit distinct differences. As shown in Figure 2 (a), adversarial attacks compute input-specific perturbations, whereas backdoor attacks apply a consistent trigger across inputs. To elaborate, injecting adversarial perturbations into a data sample forces it to shift in an input-specific direction within the representation space, ultimately crossing the decision boundary and altering its prediction. In contrast, a backdoor attack introduces a trigger pattern into each sample, effectively guiding all affected samples toward a consistent direction in the representation space, irrespective of their original features. Notably, a model compromised by a backdoor attack typically learns distinct regions corresponding to the attack target label, forming a backdoor plane in the figure within the representation space. Additionally, adversarial attacks can target any model, while backdoor attacks affect only models trained with trigger samples.
@@ 4. Threat Model and Assumptions
Assumptions. We assume that a federated learning system with numerous clients participating without directly sharing their local training data with the central server. Additionally, we assume that malicious clients comprise less than 50% of the total number of participating clients(Fang et al., 2020). We consider these assumptions reasonable for practical federated learning environments and associated attack scenarios. In addition, we assume that the server has a few clean validation samples that can be used for additional processes (Cao et al., 2020).
@@ 5. Framework Design
Figure 3. Overall workflow of DeTrigger. DeTrigger leverages insights from adversarial attack methodologies to effectively identify the trigger and prune the backdoor knowledge. The server first distributes the global model to selected clients for the local model training. Malicious clients may introduce backdoor triggers into their datasets. To mitigate this, DeTrigger analyzes gradients to detect potential backdoor triggers and prunes suspicious models before updating the global model.
@@ 5.2. Gradient Preprocessing
Using these input layer gradients, DeTrigger identifies and filters additive elements within each model’s gradient that shift predictions from the original to the target attack label. As also demonstrated in our preliminary studies (Figure 4 (c)), this process extracts information within the gradients specifically linked to the trigger pattern. Since the fundamental concept of the backdoor attack involves "injecting" a malicious trigger pattern to manipulate the model into producing an incorrect prediction, DeTrigger selectively utilizes only the elements that exhibit positive perturbations in shifting the original prediction toward the target label. As shown in Figure 4 (ii), the filtered gradient not only includes the backdoor trigger-related feature (red pixels in the center) but also input-specific features. To minimize sample-specific noise variations, DeTrigger averages the gradient elements across multiple samples used for extracting the input layer gradient. The averaged gradients are then normalized through min-max scaling, and a mask is generated to emphasize spatial locations where gradient amplitudes exceed a set threshold. As the sample depicted in Figure 4 (iii), this averaging and thresholding process allows DeTrigger to focus on regions with stronger trigger-related signals, effectively filtering out irrelevant information. In this paper, we set the threshold to 0.5.
@@
Figure 5. (a) Conceptual illustration of the impact of temperature scaling on normal data feature space and backdoor feature space. (b) Sample of ground truth and inferred triggers with different temperatures. (c) L1-norm between ground truth and inferred triggers with varying temperatures.
@@
Figures 5 (b) and (c) illustrate inferred trigger patterns and the impact of temperature scaling on the L1-norm, respectively. Specifically, the results show the L1-norm of gradients decreases with moderate temperature scaling ($T>1$), improving the clarity of inferred trigger patterns. We observe that increasing the temperature shifts the sample’s position within the normal data plane closer to the decision boundary while simultaneously reducing the influence of irrelevant features (i.e., input-specific noise), as illustrated in Figure 5 (b). This phenomenon implies that a higher temperature reduces the required perturbation magnitude ($\epsilon$ in Figure 5 (a)) to alter the decision toward the backdoor attack target. Consequently, this supports our hypothesis that the backdoor feature space resides near the decision boundary within the normal data plane. This insight aligns with prior findings by Su et al., who observed that the decision boundary of the backdoor sample is tangent to other labels (Su et al., 2024). This suggests gradients obtained with smoothed probability distributions (i.e., higher temperatures) can more effectively capture information related to the backdoor trigger. Note that we set the temperature to five in this work.